# Sufficient digits and density estimation: A Bayesian nonparametric approach using generalized finite Pólya trees

### v1

@@ 5.1.2 Posterior simulation under a GFPT2 prior
Now, consider the posterior distribution of $(Y,R,N)$ for the case (14) where a GFTP2 prior has been specified. The following theorem derives the posterior distribution of first $N\,|\,R$, second $R\,|\,N$, and third $Y\,|\,(N,R)$, where we use the following notation. Recall the definitions (12) and (19). Let $R_{0}=(R_{x_{1:n-1},0}\,|\,x_{1:n-1}\in\{0,1\}^{n-1},n\in\mathbb{N})$. For $j\in\mathbb{N}_{0}$, to stress the dependence of $R^{(j)}_{0}$ (or equivalently of $R^{(j)}$), we use capital letters: For $x_{1:j}\in\{0,1\}^{n}$, let $N_{x_{1:j}}=n_{x_{1:j}}(z_{1:m})$ be the the number of observations $z_{i}$ falling in $I_{x_{1:j}}$, which depends on $R^{(j)}$. Similarly, let $\Gamma_{x_{1:j}}=\gamma_{x_{1:j}}(z_{1:m},\alpha)$ and $L_{x_{1:j}}=\ell_{x_{1:j}}$, which both depend on $R^{(j)}$. For $n\in\mathbb{N}$, $Y^{(n)}$ can be identified by $Y^{(n)}_{0}=(Y_{x_{1:j-1},0}\,|\,x_{1:j-1}\in\{0,1\}^{j-1},j=1,\ldots,n)$ since $Y_{x_{1:j-1},1}=1-Y_{x_{1:j-1},0}$. Let $Y^{(0)}_{0}=Y^{(0)}=\emptyset$. For $n\in\mathbb{N}_{0}$, the dimension of $Y^{(n)}_{0}$ is $12^{n}-$ 1. For $j\in\mathbb{N}$ and $x_{1:j-1}\in\{0,1\}^{j-1}$, let $\nu_{x_{1:j-1}}$ be the Lebesgue measure on $[0,1)$ if $\alpha_{x_{1:j-1},0}>0$ and $\alpha_{x_{1:j-1},1}>0$, and $\nu_{x_{1:j-1}}$ be the Dirac measure concentrated at $k\in\{0,1\}$ if $\alpha_{x_{1:j-1},k}=0$ and $\alpha_{x_{1:j-1},1-k}>0$. For $n\in\mathbb{N}$, let $A_{n}=[0,1)^{2^{n}-1}$ be equipped with the corresponding Borel $\sigma$-algebra $\mathcal{F}_{n}$. Let $\mu_{n}$ be the product measure on $\mathcal{F}_{n}$ given by
@@
$p(n\,|\,R,z_{1:m})\propto p_{n}\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{b}\big{(}\Gamma_{x_{1:j-1},0},\Gamma_{x_{1:j-1},1}\big{)}}{\prod_{x_{1:n}\in\{0,1\}^{n}}L_{x_{1:n}}^{N_{x_{1:n}}}}$ (25)
@@
$p(r_{0}^{(n)}\,|\,n,z_{1:m})\propto\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{b}\big{(}\Gamma_{x_{1:j-1},0},\Gamma_{x_{1:j-1},1}\big{)}r_{x_{1:j-1,0}}^{\beta_{x_{1:j-1,0}}-1}r_{x_{1:j-1,1}}^{\beta_{x_{1:j-1,1}}-1}}{\prod_{x_{1:n}\in\{0,1\}^{n}}L_{x_{1:n}}^{N_{x_{1:n}}}}$ (27)
@@
5.2 Consistency for the posterior of $N$
In this section, we consider the posterior distribution for the case (10) where a GFTP1 prior has been specified but assume $Z_{1},\ldots,Z_{m}$ are i.i.d. according to a PDF $f^{*}$, which we refer to as the true distribution of $Z_{1:m}$. We focus on the asymptotic behaviour of the latent variable $N$ when $m\rightarrow\infty$. At the end of Section 5.2.1 (Remark 15) we motivate the use of GFPT2 priors.
@@ 5.2 Consistency for the posterior of $N$
The last inequality in (30) is akin of local stability condition, cf. [30].
@@ 5.2.1 When $N$ is bounded or not
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}p(n\,|\,Z_{1:m})>0\right){=0.}$ (33)
@@ 5.2.2 Round-off errors
We now analyse the impact on posterior inference caused by two types of round-off errors in the data.
@@
Theorem 5.6 shows that the round-off error introduced by using 0’s after the precision $\bar{n}$ leads to an inconsistent posterior distribution of $N$ as $m\to\infty$. This is still true if we use another fixed digits rule than in (34), e.g., if we replace the 0’s by 1’s after level $\bar{n}$.
The following corollary follows immediately from Theorems 3.2 and 5.4. For $i=1,\ldots,m$, let $Z_{i}=.X^{i}_{1}X^{i}_{2}\ldots$.
@@
.$X^{i}_{\bar{n}+1}X^{i}_{\bar{n}+2}\ldots\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\mathrm{Unif}[0,1)\quad\mbox{for }i=1,\ldots,m.$ (35)
@@
Under the true distribution, if $n^{*}$ digits are sufficient, then $n^{*}$ is the largest possible value of $N$. Ideally we should therefore have that $n^{*}\in\mathcal{N}\subseteq\{0,\ldots,\bar{n}\}$ but in practice we do not know the value of $n^{*}$ (provided it exists). This suggests to let $\mathcal{N}=\{0,\ldots,\bar{n}\}$. This is also an intuitive suggestion as the data contains no information beyond level $n^{*}$ of the nested binary partitions.
@@ 5.3 Consistency for the posterior of $f$
In this section we assume the same setting as in Section 5.2 and consider the random GFPT1 PDF $f(\cdot\,|\,Y^{(N)})$ and the random GFPT2 PDF $f(\cdot\,|\,Y^{(N)},R^{(N)})$ given by (6) and (13), respectively. For $i=1,2$, using the short hand notation $f=f(\cdot\,|\,Y^{(N)})$ if $i=1$ and $f=f(\cdot\,|\,Y^{(N)},R^{(N)})$ if $i=2$, recall that a posteriori $f\,|\,Z_{1:m}\sim\Pi_{i}(\cdot\,|\,Z_{1:m})$, cf. (11) and (15). As the sample size increases, we will show that $\Pi_{i}(\cdot\,|\,Z_{1:m})$ concentrates on infinitesimal small neighbourhoods of $f^{*}$ with respect to the Kolmogorov-Smirnov distance $d_{\mathrm{KS}}$ and the Hellinger distance $d_{\mathrm{H}}$: For two PDFs $f_{1}$ and $f_{2}$ on $[0,1)$, if $F_{i}(x)=\int_{0}^{x}f_{i}(x)\,\mathrm{d}x$ for $i=1,2$ and $10\leq x<$ 1, recall that
@@
Theorem 5.8 gives sufficient conditions for the posterior of $f$ to concentrate on vanishing neighbourhoods of $f^{*}$ in the weak topology, since convergence with respect to $d_{\mathrm{KS}}$ is stronger than weak covergence of probability measures on Euclidean spaces (see e.g. Theorem 6 in [15]). However, as discussed in [14], consistency under the weak topology is often considered “too weak” for density estimation purposes. Instead the next theorem gives sufficient conditions for consistency with respect to the Hellinger distance. Convergence with respect to $d_{\mathrm{H}}$ is stronger than weak convergence, and the topology induced by the Hellinger distance is equivalent to the one induced by the total variation norm (see e.g. [24]).
@@
For instance, the condition $0<p_{n}<Ce^{-ck}$ is satisfied by the Poisson distribution. By contrast, posterior consistency for standard Pólya tree priors was established in [2] (see also Theorem 7.16 in [14]) under the assumptions that the parameters $\alpha_{x_{1:n}}$ grow exponentially fast with $n$. As noted in [14] this assumption is too strong, since practitioners usually follow [25] by setting $\alpha_{x_{1:n}}=\alpha_{0}n^{2}$. See also [7, 16] for further developments.
6 Modelling multi-scale human data via Benford’s law and generalized Pólya trees
Real-world datasets – such as those involving incomes, city sizes, or seismic magnitudes – often span multiple orders of magnitude [8]. Modelling such data poses a challenge for Bayesian density estimation because traditional methods may have difficulty capturing both the coarse-scale variability and the fine-scale structure [35]. We show here how a simple extension of a GFPT1 prior can be used to obtain a simple, yet powerful, Bayesian model for density estimation in scenarios where data span multiple orders of magnitudes and their digits exhibit the Newcomb–Benford law. The main idea lies in independently modelling the order of magnitude of the data and their digits, by assuming a GFPT1 prior for the digits.
[…]
$\mathrm{P}(M_{i}=k\,|\,\omega)=\omega_{k},\quad k\in T,$
where we impose the prior
$\omega=(\omega_{k}\,|\,k\in T)\sim\mathrm{Dir}(\eta)$
The $\mathrm{MBPT}_{q}$ prior induces a scale invariant (in the terminology of [22]) random density on $T\times[0,1)$ equipped with the product of counting measure on $T$ and Lebesgue measure on $[0,1)$. This random density is given by
Posterior computation is straightforward. Indeed, $(Y,N)$ and $\omega$ are independent also a posteriori, with the posterior of $(Y,N)\,|\,Z_{1:m}$ given by a slight modification of Theorem 5.1. Moreover, $\omega\,|\,Z_{1:m}\sim\mathrm{Dir}(\eta^{p})$ where $\eta^{p}_{k}=\eta_{k}+\sum_{i=1}^{m}I[M_{i}=k]$ for $k\in T$.
@@ 7.1 Posterior results when using GFPT1 and Pólya tree priors
[…]
Figure 1 shows the posterior density estimates (see Remark 12) for the 12 cases. For all datasets, the PT prior yields extremely wiggly posterior density estimates when $\alpha_{0}=0.05$ (dotted orange line), and these get only partially less wiggly for $\alpha_{0}=2$ (dotted blue line). On the other hand, for datasets DG1 and DG2 (the two first columns), the GFPT1 prior yields precise posterior density estimates for both values of $\alpha_{0}$ (the solid blue and orange lines). However, for DG3 and $\alpha_{0}=2.0$, the posterior density estimate under the GFPT1 prior is not very precise (last column, solid blue line). This happens because the posterior of $N$ concentrates on small values leading to a coarse approximation. When $\alpha_{0}=0.05$ instead, the posterior of $N$ concentrates to higher values leading to a more accurate density estimate (last column, solid orange line).
@@
    the total variation distance distance between the density of the true data generating process and the posterior density estimate (for short the true and estimated densities),
@@
respectively. As expected Figures 24 show that larger values of $\alpha_{0}$ correspond to smaller values of $N$ (a posteriori) and vice versa. This is particularly evident for datasets DG1 and DG3. In general, we see that GFPT1 priors outperform PT priors in terms of the quality of the posterior density estimate. Moreover, larger $\alpha_{0}$ clearly yield less wiggly functions when using PT priors, while this dependence is not so clear when using GFPT1 priors. In conclusion, we find that $\alpha_{0}=0.1$ provides the best trade off between wigglyness and accuracy of the density estimate, and we suggest using it as default value for GFPT1 priors.

### v2

@@ 5.1.2 Posterior simulation under a GFPT2 prior
Now, consider the posterior distribution of $(Y,R,N)$ for the case (14) where a GFPT2 prior has been specified. The following theorem derives the posterior distribution of first N|RN\,|\,R, second R|NR\,|\,N, and third Y|(N,R)Y\,|\,(N,R), where we use the following notation. Recall the definitions (12) and (19). Let $R_{0}=(R_{x_{1:n-1},0}\,|\,x_{1:n-1}\in\{0,1\}^{n-1},n\in\mathbb{N})$. For $j\in\mathbb{N}_{0}$, to stress the dependence of $R^{(j)}_{0}$ (or equivalently of $R^{(j)}$), we use capital letters: For $x_{1:j}\in\{0,1\}^{n}$, let $N_{x_{1:j}}=n_{x_{1:j}}(z_{1:m})$ be the the number of observations $z_{i}$ falling in $I_{x_{1:j}}$, which depends on $R^{(j)}$. Similarly, let $\Gamma_{x_{1:j}}=\gamma_{x_{1:j}}(z_{1:m},\alpha)$ and $L_{x_{1:j}}=\ell_{x_{1:j}}$, which both depend on $R^{(j)}$. For $n\in\mathbb{N}$, $Y^{(n)}$ can be identified by $Y^{(n)}_{0}=(Y_{x_{1:j-1},0}\,|\,x_{1:j-1}\in\{0,1\}^{j-1},j=1,\ldots,n)$ since $Y_{x_{1:j-1},1}=1-Y_{x_{1:j-1},0}$. Let $Y^{(0)}_{0}=Y^{(0)}=\emptyset$. For $n\in\mathbb{N}_{0}$, the dimension of $Y^{(n)}_{0}$ is $2^{n}-1$. For $j\in\mathbb{N}$ and $x_{1:j-1}\in\{0,1\}^{j-1}$, let $\nu_{x_{1:j-1}}$ be the Lebesgue measure on $[0,1)$ if $\alpha_{x_{1:j-1},0}>0$ and $\alpha_{x_{1:j-1},1}>0$, and $\nu_{x_{1:j-1}}$ be the Dirac measure concentrated at $k\in\{0,1\}$ if $\alpha_{x_{1:j-1},k}=0$ and $\alpha_{x_{1:j-1},1-k}>0$. For $n\in\mathbb{N}$, let $A_{n}=[0,1)^{2^{n}-1}$ be equipped with the corresponding Borel $\sigma$-algebra $\mathcal{F}_{n}$. Let $\mu_{n}$ be the product measure on $\mathcal{F}_{n}$ given by
@@
$p(n\,|\,R,z_{1:m})\propto p_{n}\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{beta}\big(\Gamma_{x_{1:j-1},0},\Gamma_{x_{1:j-1},1}\big)}{\prod_{x_{1:n}\in\{0,1\}^{n}}L_{x_{1:n}}^{N_{x_{1:n}}}}$ (25)
@@
$p(r_{0}^{(n)}\,|\,n,z_{1:m})\propto\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{beta}\big(\Gamma_{x_{1:j-1},0},\Gamma_{x_{1:j-1},1}\big)r_{x_{1:j-1,0}}^{\beta_{x_{1:j-1,0}}-1}r_{x_{1:j-1,1}}^{\beta_{x_{1:j-1,1}}-1}}{\prod_{x_{1:n}\in\{0,1\}^{n}}L_{x_{1:n}}^{N_{x_{1:n}}}}$ (27)
@@
5.1.3 Posterior simulation under a MBPT prior
Posterior computation is straightforward under a MBPT prior as given in Section $4.$ Indeed, $(Y,N)$ and $\omega$ are independent also a posteriori, with the posterior of $(Y,N)\,|\,Z_{1:m}$ given by a slight modification of Theorem 5.1. Moreover, $\omega\,|\,Z_{1:m}\sim\mathrm{Dir}(\eta^{p})$ where $\eta^{p}_{k}=\eta_{k}+\sum_{i=1}^{m}I[M_{i}=k]$ for $k\in T$.
5.2 Consistency for the posterior of NN
In this section, we consider the posterior distribution for the case (10) where a GFPT1 prior has been specified but assume $Z_{1},\ldots,Z_{m}$ are i.i.d. according to a PDF $f^{*}$, which we refer to as the true distribution of $Z_{1:m}$. We focus on the asymptotic behaviour of the latent variable $N$ when $m\rightarrow\infty$. At the end of Section 5.2.1 (Remark 16) we motivate the use of GFPT2 priors.
@@ 5.2 Consistency for the posterior of NN
The last inequality in (30) is akin of local stability condition, cf. Møller and Waagepetersen (2004).
@@ 5.2.1 When $N$ is bounded or not
$\mathrm{P}^{*}\left(\lim_{m\rightarrow\infty}p(n\,|\,Z_{1:m})>0\right){\color[rgb]{0,0,0}=0.}$ (33)
@@ 5.2.2 Round-off errors
We now analyse the impact on posterior inference caused by two types of round-off errors in the data. Specifically, we assume that data are observed with a precision of $\bar{n}\in\mathbb{N}_{0}$ digits and consider two scenarios. In the first one, we ignore the round-off and assume to have observed the data perfectly: We show in Theorem 5.6 that this leads to an inconsistent posterior for $N$ as $m\rightarrow\infty$ even if $n^{*}\leq\bar{n}$. In the second scenario, we explicitly account for the round-off and acknowledge that data contain no information beyond digit $\bar{n}$. That is, we assume that we do not actually observe an i.i.d. sample from $f^{*}$ but rather from $\bar{f}^{*}$, which is a piecewise constant approximation of $f^{*}$ at level $\bar{n}$ of the nested binary partition. In such a case, if $n^{*}\leq\bar{n}$, we obtain posterior consistency for $N$ in Corollary 5.7.
@@
The conclusion of Theorem 5.6 is unchanged if we use another fixed digits rule than in (34), e.g., if we replace the 0’s by 1’s after level $\bar{n}$.
We now consider the scenario where we explicitly account for the round-off error. The following corollary follows immediately from Theorems 3.2 and 5.4. For $i=1,\ldots,m$, let $Z_{i}=.X^{i}_{1}X^{i}_{2}\ldots$.
@@
$X^{i}_{\bar{n}+1}X^{i}_{\bar{n}+2}\ldots\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\mathrm{Unif}[0,1)\quad\mbox{for }i=1,\ldots,m.$ (35)
@@
Under the true distribution $f^{*}$, if $n^{*}$ digits are sufficient, then $n^{*}$ is the largest possible value of $N$. Ideally we should therefore have that $n^{*}\in\mathcal{N}\subseteq\{0,\ldots,\bar{n}\}$ but in practice we do not know the value of $n^{*}$ (provided it exists). This suggests to let $\mathcal{N}=\{0,\ldots,\bar{n}\}$. This is also an intuitive suggestion as the data contains no information beyond level $n^{*}$ of the nested binary partitions.
On the other hand, if under the true distribution $n^{*}>\bar{n}$, then the posterior of $N$ will concentrate on $\bar{n}$. This is intuitive as assuming (35) means that we do not observe an i.i.d. sample from $f^{*}$ but from a density $\bar{f}^{*}$ which is piecewise constant beyond level $\bar{n}$ of the nested binary partition.
@@ 5.3 Consistency for the posterior of $f$
In this section we assume the same setting as in Section 5.2 and consider the random GFPT1 PDF $f(\cdot\,|\,Y^{(N)})$ and the random GFPT2 PDF $f(\cdot\,|\,Y^{(N)},R^{(N)})$ given by (6) and (13), respectively. For $i=1,2$, using the short hand notation $f=f(\cdot\,|\,Y^{(N)})$ if $i=1$ and $f=f(\cdot\,|\,Y^{(N)},R^{(N)})$ if $i=2$, recall that a posteriori $f\,|\,Z_{1:m}\sim\Pi_{i}(\cdot\,|\,Z_{1:m})$, cf. (11) and (15). As the sample size increases, we will show that $\Pi_{i}(\cdot\,|\,Z_{1:m})$ concentrates on infinitesimal small neighbourhoods of $f^{*}$ with respect to the Kolmogorov-Smirnov distance $d_{\mathrm{KS}}$ and the Hellinger distance $d_{\mathrm{H}}$: For two PDFs $f_{1}$ and $f_{2}$ on $[0,1)$, if $F_{i}(x)=\int_{0}^{x}f_{i}(x)\,\mathrm{d}x$ for $i=1,2$ and $0\leq x<1$, recall that
@@
Theorem 5.8 gives sufficient conditions for the posterior of $f$ to concentrate on vanishing neighbourhoods of $f^{*}$ in the weak topology, since convergence with respect to $d_{\mathrm{KS}}$ is stronger than weak covergence of probability measures on Euclidean spaces (see e.g. Theorem 6 in Gibbs and Su, 2002). However, as discussed in Ghoshal and Vaart (2017), consistency under the weak topology is often considered “too weak” for density estimation purposes. Instead the next theorem gives sufficient conditions for consistency with respect to the Hellinger distance. Convergence with respect to $d_{\mathrm{H}}$ is stronger than weak convergence, and the topology induced by the Hellinger distance is equivalent to the one induced by the total variation norm (see e.g. Kraft, 1955).
@@
For instance, the condition $0<p_{n}<Ce^{-ck}$ is satisfied by the Poisson distribution. By contrast, posterior consistency for standard Pólya tree priors was established in Barron et al. (1999) (see also Theorem 7.16 in Ghoshal and Vaart, 2017) under the assumptions that the parameters $\alpha_{x_{1:n}}$ grow exponentially fast with $n$. As noted in Ghoshal and Vaart (2017) this assumption is too strong, since practitioners usually follow Lavine (1992) by setting $\alpha_{x_{1:n}}=\alpha_{0}n^{2}$. See also Castillo (2017) and Giordano et al. (2025) for further developments.
@@ 6.1 Posterior results when using GFPT1 and Pólya tree priors
where $\mathcal{N}(\mu,\sigma^{2})|_{[0,1)}$ denotes the truncation on the unit interval of the normal distribution with mean $\mu$ and variance $\sigma^{2}$. We simulated $m=1000$ observations from each data generating process and fitted them using either a $\mathrm{GFPT1}(\alpha,I)$ or a $\mathrm{PT}(\alpha,I)$ prior where $\alpha$ is given by (16) with $\alpha_{0}=2$ or $\alpha_{0}=0.05$ and where $I$ is the sequence of standard diadic partitions. Thus, in our first simulation study, we consider 12 cases corresponding to the four different models for each of the three datasets.
Figure 1 shows the density estimates (as given in Remark 13) for the 12 cases. For all datasets, the PT prior yields extremely wiggly density estimates when $\alpha_{0}=0.05$ (dotted orange line), and these get only partially less wiggly for $\alpha_{0}=2$ (dotted blue line). On the other hand, for datasets DG1 and DG2 (the two first columns), the GFPT1 prior yields precise density estimates for both values of $\alpha_{0}$ (the solid blue and orange lines). However, for DG3 and $\alpha_{0}=2.0$, the density estimate under the GFPT1 prior is not very precise (last column, solid blue line). This happens because the posterior of $N$ concentrates on small values leading to a coarse approximation. When $\alpha_{0}=0.05$ instead, the posterior of $N$ concentrates to higher values leading to a more accurate density estimate (last column, solid orange line).
Next, we focus on the impact that assuming $N$ random has on prior and posterior variability. For the GFPT1 prior, we assume that a priori $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,20\}$ where the Poisson distribution has mean 10. As alternative, we consider a finite PT model which corresponds to setting $N=\delta_{10}$. For both models, $\alpha$ is given by (16) for $\alpha_{0}\in\{0.05,0.1,2.0,10.0\}$. Figure 2 shows global credible bands for the random PDF under both priors: A priori, the introduction of a prior on $N$ does not seem to make a huge difference in terms of variability. However, the posterior under the two models is strikingly different, as shown in Figure 3: For all choices of the parameter $\alpha_{0}$ and under all three data generating processes (DG1)–(DG3), the posterior under the GFPT1 is much more concentrated, highlighting the benefits of assuming a prior for $N$ as opposed to fixing it to a large value. Moreover, Figure 3 (bottom row) shows the effect that $\alpha_{0}$ has on the posterior of $N$: Large values of $\alpha_{0}$ shrink $N$ to smaller values a posteriori, resulting in coarser density estimates. In summary, under a GFPT1 prior, we find that specifying a small value for $\alpha_{0}$ enables greater adaptability via the prior on $N$ that regularize the density estimate.
Figure 2: 95% global credible bands (shared area) and mean (solid line) of the random PDF under a GFPT1 prior or a finite PT prior when $\alpha$ is given by (16) for $\alpha_{0}\in\{0.05,0.1,2.0,10.0\}$. For the GFPT1 prior, a priori $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,20\}$ where the Poisson distribution has mean 10. For the finite PT prior, the NBP $I$ has depth 10.
Figure 3: 95% global credible bands (shared area) and mean (solid line) of the posterior of the random PDF under the same specifications of Figure 2, when data are generated by (DG1)–(DG3) (different rows, from top to bottom)
@@
    the total variation distance distance between the density of the true data generating process (for short the true density) and the density estimate (cf. Remark 13),
@@
respectively. As expected Figures 46 show that larger values of $\alpha_{0}$ correspond to smaller values of $N$ (a posteriori) and vice versa. This is particularly evident for datasets DG1 and DG3. In general, we see that GFPT1 priors outperform PT priors in terms of the quality of the density estimate. Moreover, larger $\alpha_{0}$ clearly yield less wiggly functions when using PT priors, while this dependence is not so clear when using GFPT1 priors. In conclusion, we find that $\alpha_{0}=0.1$ provides the best trade off between wigglyness and accuracy of the density estimate, and we suggest using it as default value for GFPT1 priors.